# Gravitational Waves from Type-I Strings in a Neutrino Mass Model

### v1

@@ III Type-I Cosmic Strings
A key feature of our framework is the natural emergence of type-I CSs, characterized by $\beta=\lambda/(2g^{2})<1$. This is achieved through a small quartic coupling $\lambda$, which is well-motivated within the split seesaw mechanism. In conventional 44D seesaw models, the small neutrino Yukawa coupling often requires fine-tuning. In our higher-dimensional setup, the small effective Yukawa is generated geometrically by wavefunction suppression, Eq. 13. This liberates the quartic coupling $\lambda$ from the obligation to be large. In fact, a small value of $\lambda$ is technically natural. This naturalness is explicitly realized in supersymmetric (SUSY) extensions. In SUSY theories, the quartic couplings of scalar fields are not free parameters but are determined by the gauge couplings, leading to a natural expectation of $\lambda\sim O(0.1)$ [28]. While our present construction is non-SUSY, it is motivated by the fact that a small $\lambda$ is a robust feature of well-motivated UV completions. This distinguishes our scenario from a finely-tuned one and provides a strong theoretical motivation for exploring the $\beta<1$ regime.
@@
Furthermore, the value of $\beta$ profoundly impacts the CS network evolution and its GW signature. While the expression for $\Omega_{\text{GW}}(f)$ in Eq. 18 has a generic form, it depends critically on the string tension $\mu\propto\nu_{R}^{2}\,B(\beta)$. The function $B(\beta)$ varies from $1$ for $\beta\ll 1$ to $ln(\beta)$ for $\beta\gg 1$, leading to a significantly different normalization of the GW spectrum for the same symmetry-breaking scale $\nu_{R}$. More importantly, the parameter $\beta$ determines the fraction of energy lost to gravitational radiation versus particle emission, which influences the loop number density (parameterized by $\zeta$ and $\alpha$ in Eq. 18) [29, 30]. Type-I strings have attractive self-interactions, leading to a different loop-chopping efficiency and network scaling solution compared to the more commonly studied type-II case. Our analysis, which focuses on the $\beta\ll 1$ regime, therefore explores a distinct and observationally viable corner of the CS parameter space that is intimately linked to the neutrino mass generation mechanism. Our analysis is now extended to the case $\beta\simeq 0.01$.
@@ III.1 Phenomenological Implications
Let us now explain the GW power spectrum for type-I cosmic strings. We follow the treatment given in [10]. The total GW density at the present frequency $f$ is,
$\Omega_{\rm GW}(f)$ $\simeq\dfrac{2\,\Gamma\,G\mu^{2}}{\rho_{c}\,f}\,\dfrac{\zeta}{\alpha\left(\alpha+\Gamma\,G\mu\right)}\times\int_{t_{f}}^{t_{0}}d\tilde{t}\left(\dfrac{a(\tilde{t})}{a(t_{0})}\right)^{5}$ (18)
@@
Here $\Gamma\simeq 50$ quantifies the efficiency of GW emission, $G\mu$ is the dimensionless string tension, $\rho_{c}=3H_{0}^{2}/(8\pi G)$ is the critical density ($H_{0}$ is the Hubble parameter today and $G$ is Newton’s constant). We considered the largest width of the loop parameterized by $\alpha=0.1$ and $\zeta=10$, which quantifies the net energy flux associated with the loops. The upper limit on the integration is up to the present time $t_{0}$, the lower limit $t_{f}$ is the time when the cosmic strings enter the scaling regime, $a(t)$ being the scale factor and $t_{i}$ is the loop formation time given as,
$t_{i}=\dfrac{\tilde{l}+\Gamma\,G\mu\,\tilde{t}}{\alpha+\Gamma\,G\mu}.$ (19)
Here $\tilde{l}=2a(\tilde{t})/(fa(t_{0}))$ is the size of the loop. We have presented the GW power spectra in Fig. 2, which indicates the testability of the model and the symmetry-breaking scale in the present and future planned experiments. The different colors in the power spectrum indicate the benchmark point for the symmetry-breaking scale as indicated in Fig. 1. To be consistent with the current PTA dataset [27], the symmetry-breaking scale $\nu_{R}\lesssim 3\times 10^{13}$ GeV as shown by the gray dashed line in Fig. 2. Therefore, the emission of GWs from type-I CSs provides testable signatures of neutrino masses through cosmology.
Figure 2: The evolution of GW spectra for type-I CSs for different values of $\nu_{R}$ corresponding to the benchmark points given in Fig. 1 and indicated in the inset. The dashed gray line corresponds to $\nu_{R}\simeq 3\times 10^{13}$, consistent with the PTAs. The colored shaded regions indicate the sensitivity curves of present (solid boundaries) LIGO O3 [31], NANOGrav [27] and future (dashed boundaries) LIGO O5, SKA [32], THEIA [33], LISA [34], $\mu$-ARES [35], BBO [36], U-DECIGO [37, 38], CE [39] and ET [40] experiments.
@@ IV Conclusions
This work adopts the left-right symmetric model that provides a natural gauge-invariant embedding of right-handed neutrinos as part of $SU(2)_{R}$ doublets. By incorporating the split seesaw mechanism into this framework, one achieves exponentially suppressed effective Yukawa couplings due to wavefunction localization in extra dimensions, while the heavy Majorana mass $M_{N}$ arises from the spontaneous breaking of $SU(2)_{R}\times U(1)_{B-L}$. This unified picture not only explains the smallness of neutrino masses without fine-tuning but also predicts the formation of type-I cosmic strings with $\beta\ll 1$ associated with the BLB–L breaking scale. In the split seesaw mechanism, one can explain the matter-antimatter asymmetry of the universe, as well as dark matter, which we leave for future work, but some discussion has already been made in Ref. [26]. This framework, therefore, simultaneously addresses neutrino physics and early-universe cosmology in a consistent high-energy theory.
@@ Acknowledgments
The author gratefully acknowledges Qaisar Shafi, Rishav Roshan, Rinku Maji, and Anish Ghoshal for their valuable discussions.

### v2

@@ III Type-I Cosmic Strings
The parameter $\beta=\lambda/(2g^{2})$ critically determines the nature of the CSs. For $\beta<1$, the strings are of type-I and exhibit attractive self-interactions, in contrast to the repulsive interactions of type-II strings ($\beta>1$). As a consequence of this attraction, higher winding number bound states can be formed. In particular, these bound states prevent an ordered Abrikosov lattice from existing in laboratory experiments involving type I superconductors [10]. We refer the reader to Ref. [35] for detailed analysis on the evolution of CS loops and to Refs. [43, 39] for detailed analysis on type-I and II CSs. A key feature of our framework is the natural emergence of type-I CSs, characterized by $\beta<1$. This is achieved through a small quartic coupling $\lambda$, which is well-motivated within the split seesaw mechanism. In conventional 44D seesaw models, the small neutrino Yukawa coupling often requires fine-tuning. In our higher-dimensional setup, the small effective Yukawa is generated geometrically by wavefunction suppression, Eq. 22. This liberates the quartic coupling $\lambda$ from the obligation to be large. In fact, a small value of $\lambda$ is technically natural. This naturalness is explicitly realized in supersymmetric (SUSY) extensions. In SUSY theories, the quartic couplings of scalar fields are not free parameters but are determined by the gauge couplings, leading to a natural expectation of $\lambda\sim O(0.1)$ [27]. While our present construction is non-SUSY, it is motivated by the fact that a small $\lambda$ is a robust feature of well-motivated UV completions. This distinguishes our scenario from a finely-tuned one and provides a strong theoretical motivation for exploring the $\beta<1$ regime.
@@
Furthermore, the value of $\beta$ profoundly impacts the CS network evolution and its GW signature. While the expression for $\Omega_{\text{GW}}(f)$ in Eq. 29 has a generic form, it depends critically on the string tension $\mu_{cs}\propto v_{R}^{2}\,B(\beta)$. The function $B(\beta)$ varies from $1$ for $\beta\ll 1$ to $ln(\beta)$ for $\beta\gg 1$, leading to a significantly different normalization of the GW spectrum for the same symmetry-breaking scale $v_{R}$. More importantly, the parameter $\beta$ determines the fraction of energy lost to gravitational radiation versus particle emission, which influences the loop number density (parameterized by $\zeta$ and $\alpha$ in Eq. 29) [35, 12]. Type-I strings have attractive self-interactions, leading to a different loop-chopping efficiency and network scaling solution compared to the more commonly studied type-II case. Our analysis, which focuses on the $\beta\ll 1$ regime, therefore explores a distinct and observationally viable corner of the CS parameter space that is intimately linked to the neutrino mass generation mechanism. We have shown the impact of $\beta$ on the GW spectrum in Fig. 2 for a fixed value of $v_{R}=10^{14}$ GeV. The GW spectrum decreases logarithmically with the decrease in $\beta$.
Figure 2: The evolution of GW spectra from CSs for different values of $\beta$ as shown by the vertical colored bar. We fix $v_{R}=10^{14}$ GeV, the amplitude of the GW spectrum decreases logarithmically with the decrease in $\beta$.
Our analysis is now extended to the case $\beta\simeq 0.01$ to put the constraints on the parameter space for a wide range of GW detectors.
Figure 3: The evolution of GW spectra for type-I CSs for different values of $v_{R}$ corresponding to the benchmark points given in Fig. 1 and indicated in the inset. The dashed gray line corresponds to $v_{R}\simeq 1\times 10^{15}$, which is not consistent with the PTAs and hence ruled out. The colored shaded regions indicate the sensitivity curves of present (solid boundaries) LIGO O3 [1], NANOGrav [4] and future (dashed boundaries) LIGO O5, SKA [40], THEIA [17], LISA [9], $\mu$-ARES [38], BBO [14], U-DECIGO [42, 21], CE [34] and ET [32] experiments.
@@ III.1 Phenomenological Implications
Let us now explain the GW power spectrum for type-I CSs. We follow the treatment given in [15]. The total GW density at the present frequency $f$ is,
$\Omega_{\rm GW}(f)$ $\simeq\dfrac{2\,\Gamma\,G\mu_{cs}^{2}}{\rho_{c}\,f}\,\dfrac{\zeta}{\alpha\left(\alpha+\Gamma\,G\mu_{cs}\right)}\times\int_{t_{f}}^{t_{0}}d\tilde{t}\left(\dfrac{a(\tilde{t})}{a(t_{0})}\right)^{5}$ (29)
@@
Here $\Gamma\simeq 50$ quantifies the efficiency of GW emission, $G\mu_{cs}$ is the dimensionless string tension, $\rho_{c}=3H_{0}^{2}/(8\pi G)$ is the critical density ($H_{0}$ is the Hubble parameter today and $G$ is Newton’s constant). We considered the largest width of the loop parameterized by $\alpha=0.1$ and $\zeta=10$, which quantifies the net energy flux associated with the loops. The upper limit on the integration is up to the present time $t_{0}$, the lower limit $t_{f}$ is the time when the cosmic strings enter the scaling regime, $a(t)$ being the scale factor and $t_{i}$ is the loop formation time given as,
$t_{i}=\dfrac{\tilde{l}+\Gamma\,G\mu_{cs}\,\tilde{t}}{\alpha+\Gamma\,G\mu_{cs}}.$ (30)
Here $\tilde{l}=2a(\tilde{t})/(fa(t_{0}))$ is the size of the loop. We have presented the GW power spectra in Fig. 3, which indicates the testability of the model and the symmetry-breaking scale in the present and future planned experiments. The different colors in the power spectrum indicate the benchmark point for the symmetry-breaking scale as indicated in Fig. 1. To be consistent with the current PTA dataset [4], the symmetry-breaking scale $v_{R}\lesssim 10^{15}$ GeV as shown by the gray dashed line in Fig. 3. Therefore, the emission of GWs from type-I CSs provides testable signatures of neutrino masses through cosmology.
Figure 4: The evolution of the CS tension with $\beta$ and its impact on the exponential parameter appearing in Eq. 26 is shown with the vertical colored bar. The hatch region indicates the NANOGrav [4] excluded regime. The mass of the neutrino is assumed to be the atmospheric scale mass, $M_{\nu}\simeq 0.05$ eV.
It is worth noting that in the split seesaw mechanism, though the smallness of neutrino mass is gained by the exponential suppression of Yukawa couplings as given in Eq. 26, since the breaking scale $v_{R}$ is observationally constrained for a given value of $\beta$ put an upper bound on the suppression factor as shown in the Fig. 1. The summary of the constrained parameters are shown in Fig. 4. Where the evolution of the CS tension with beta has been presented and its impact on the exponential suppression factor appearing in Eq. 26 is shown with the vertical colored bar. The hatch-shaded region shows the NANOGrav excluded region. The LIGO O3 bound are less stringent than PTAs and the collider bound correspond to the very small value of $G\mu\sim 10^{-28}$.
@@ IV Conclusions
This work adopts the left-right symmetric model that provides a natural gauge-invariant embedding of right-handed neutrinos as part of $SU(2)_{R}$ doublets. By incorporating the split seesaw mechanism into this framework, one achieves exponentially suppressed effective Yukawa couplings due to wavefunction localization in extra dimensions, while the heavy Majorana mass $M_{N}$ arises from the spontaneous breaking of $SU(2)_{R}\times U(1)_{B-L}$. This unified picture not only explains the smallness of neutrino masses without fine-tuning but also predicts the formation of type-I cosmic strings with $\beta\ll 1$ associated with the BLB–L breaking scale. We have particularly shown that the suppression factor in the split seesaw mechanism is constrained by the choice of $\beta$ giving rise to the type-1 CSs. This framework addresses the neutrino mass generation from the split seesaw and its dependence on GWs emitted by type-1 CSs. In the split seesaw mechanism, one can explain the matter-antimatter asymmetry of the universe, as well as dark matter, which we leave for future work, but some discussion has already been made in Ref. [25].
@@ Acknowledgments
The author gratefully acknowledges Qaisar Shafi, Rishav Roshan, Rinku Maji, and Anish Ghoshal for their valuable discussions regarding CSs. Special thanks are also due to the anonymous referees for their insightful comments and suggestions, which significantly improved the manuscript.